YOUR TITLE HERE

by

Your Name Here

\underline{An Abstract}

Of a thesis submitted in partial fulfillment of the

requirements for the Doctor of Philosophy degree in Physics in the Graduate College of The

University of Iowa

May 2010

ABSTRACT

This is the external abstract that you need to print, get signed, and hand in to the graduate college with your abstract title page and signature page. I believe that the limit on this is two double spaced pages, including the signature

Abstract Approved: Thesis Supervisor Title and Department Date

by

Your Name Here

A thesis submitted in partial fulfillment of the

requirements for the Doctor of Philosophy degree in Physics in the Graduate College of The

University of Iowa

May 2010

Copyright by YOUR NAME HERE

1998 All Rights Reserved

Graduate College The University of Iowa

Iowa City, Iowa

CERTIFICATE OF APPROVAL

PH.D. THESIS

This is to certify that the Ph.D. thesis of

Your Name Here

has been approved by the Examining Committee for the thesis requirement for the Doctor of Philosophy degree in Physics at the May 2010 graduation.

Thesis Committee: Your members, Thesis Supervisor

member

member

member

member This is optional; start with the word “To”; you do not need to end with a period

ACKNOWLEDGMENTS

This is also optional. Use complete sentences here.

ABSTRACT

This is the abstract that is bound with the thesis.

TABLE OF CONTENTS

LIST OF TABLES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v LIST OF FIGURES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vi CHAPTER

1 NAME IN CAPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Mixed Case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1.1 Mixed Case . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 ABE LINCOLN’S GETTYSBURG ADDRESS . . . . . . . . . . . . 2

2.1 The Beginning . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 APPENDIX

A TITLE IN CAPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 BIBLIOGRAPHY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LIST OF TABLES

v LIST OF FIGURES

vi

CHAPTER 1 NAME IN CAPS

1.1 Mixed Case sections look like this

1.1.1 Mixed Case subsections look like this

CHAPTER 2 ABE LINCOLN’S GETTYSBURG ADDRESS

2.1 The Beginning

Four score and seven years ago our fathers brought forth on this continent, a new nation, conceived in Liberty, and dedicated to the proposition that all men are created equal [2].

Now we are engaged in a great civil war, testing whether that nation, or any nation so conceived and so dedicated, can long endure. We are met on a great battle-field of that war. We have come to dedicate a portion of that field, as a final resting place for those who here gave their lives that that nation might live. It is altogether fitting and proper that we should do this.

APPENDIX A TITLE IN CAPS

REFERENCES

[1] See, for example, E. V. Bugaev, A. Misaki, V. A. Naumov, T. S. Sinegovskaya, S. I. Sinegovsky and N. Takahashi, Phys. Rev. D 58, 054001 (1998) [arXiv:hepph/9803488].

[2] A. Lincoln, Gettysburg Address, unpublished.